# Higher Order Dualities between Prime Ideals

### v1

@@ 4 A New Formula for the Chebotarev Density in arbitrary finite Galois extensions
We will prove two theorems in this section, leading up to a new formula for the Chebotarev Density. We will be proving the results in a setting of a finite Galois extension $L/K$ such that $G=Gal(L/K)$ is the corresponding Galois group. We fix a conjugacy class $C\subset G$ which will be instrumental in our proofs. As mentioned above, the following results can be viewed as generalizations to the works of Alladi, Dawsey, Sweeting, Woo, Johnson and the author himself. More precisely, the result that will be derived at the end of this section will prove to be a perfect generalization101010our result generalizes the choice of the Galois extension to equation (6.3) in [Se25].
Our first theorem in this section is a crucial one for our goal. It enables us to to bring in the second order duality (Theorem 2.3) and the Chebotarev Density into the picture. This theorem can be treated as the backbone of this whole process to deduce the formula.
@@
$\sum_{2\leq N(I)\leq X}Q^{2}_{C}(I)=c_{K}\cdot\frac{|C|}{|G|}\cdot X+O\left(\frac{X(\log\log X)^{2}}{\log X}\right)$
@@
$Q_{C}^{2}(I)$ $:=\#\left\{\mathfrak{p}\subset\mathcal{O}_{K}:\;\mathfrak{p}\text{ is a prime; }I\subset\mathfrak{p};\;M_{2}(I)=N(\mathfrak{p})\;\text{and}\;\left[\frac{L/K}{\mathfrak{p}}\right]=C,\;\mathfrak{p}\;\text{is unramified}\right\}$
By definition, it is clear that ideals in $\mathcal{O}_{K}$ contained in prime ideals only of a particular norm, do not contribute to the sum in the LHS. As done in the proof of Lemma 3.4, in a similar fashion, we define the set $S_{2}(X,\mathfrak{p})$ as the set of ideals $I$ in $\mathcal{O}_{K}$ such that $N(I)\leq X$ with a unique and non-repeating prime ideal of the largest norm and $M_{2}(I)=N(\mathfrak{p})$. We note here that for any ideal $I\subset\mathfrak{p}$, for some prime $\mathfrak{p}$ such that $N(I)\leq X$ and $M_{2}(I)=N(\mathfrak{p})$, $N(\mathfrak{p})\leq\sqrt{X}$. Therefore, using Lemma 3.1, we have that
$\sum_{2\leq N(I)\leq X}Q_{C}^{2}(I)=\sum_{\begin{subarray}{c}N(\mathfrak{p})\leq\sqrt{X}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}|S_{2}(X,\mathfrak{p})|+O\left(\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}\right)$ (4.1)
@@
$|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}N(\mathfrak{M})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{M})\leq N(\mathfrak{p})\end{subarray}}\sum_{\begin{subarray}{c}N(\mathfrak{p})<N(\mathfrak{q})\\ N(\mathfrak{M.pq})\leq X\end{subarray}}1=\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}}\sum_{\begin{subarray}{c}N(\mathfrak{M})\leq\frac{X}{N(\mathfrak{pq})}\\ M_{1}(\mathfrak{M})\leq N(\mathfrak{p})\end{subarray}}1$ (4.2)
@@
$|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{O}_{K}\\ \mathfrak{q}\text{ is a prime}\\ N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{pq})},N(\mathfrak{p})\right)$ (4.3)
@@
$\sum_{\begin{subarray}{c}N(\mathfrak{p})\leq Y\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}|S_{2}(X,\mathfrak{p})|\leq\sum_{N(\mathfrak{p})\leq Y}|S_{2}(X,\mathfrak{p})|\leq\sum_{N(\mathfrak{p})\leq Y}\Psi_{K,2}\left({X},\mathfrak{p}\right)\ll\frac{X\log Y}{\log X}$ (4.4)
@@
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\;\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)$ (4.5)
@@
$Y<N(\mathfrak{p}),\quad N(\mathfrak{p})<N(\mathfrak{q}),\quad N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{q})},\quad N(\mathfrak{p})\leq\sqrt{X}$
@@
$\hskip 56.9055pt+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\;\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{q})}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)$ (4.6)
To bring in the Chebotarev Density, we now interchange the inner sums of both the double sums in the RHS of (4.6) with suitable integrals. This is indeed one of the most crucial steps of the whole proof. We do the replacements as follow:
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\frac{|C|}{|G|}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\frac{|C|}{|G|}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
@@
$E_{1}$ $:=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\frac{|C|}{|G|}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right)$
$E_{2}$ $:=\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\frac{|C|}{|G|}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right)$
@@
$=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\;\sum_{N(\mathfrak{M})\leq\frac{X}{YN(\mathfrak{q})}}\left|\sum_{\begin{subarray}{c}\max(Y,M_{1}(\mathfrak{M}))\leq N(\mathfrak{p})\leq\min\left(N(\mathfrak{q}),\frac{X}{N(\mathfrak{Mq})}\right)\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}1-\frac{|C|}{|G|}\int_{\max(Y,M_{1}(\mathfrak{M}))}^{\min\left(N(\mathfrak{q}),\frac{X}{N(\mathfrak{Mq})}\right)}\frac{dt}{\log t}\right|$ (4.8)
@@
$\ll\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}$ (4.10)
Therefore, combining (4.7), (4.9) and (4.10), we have that
@@
$+$ $\frac{|C|}{|G|}\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right)$ (4.11)
@@
$+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right)$ (4.12)
Now, we will evaluate the pseudo-integro-sums using further estimates. Surprisingly, the very same idea as used above will work here too. Let us recall the definition of $Q^{2}(I)$:
$Q^{2}(I)$ $:=\#\{\mathfrak{p}\subset\mathcal{O}_{K}:\;I\subset\mathfrak{p};\;M_{2}(I)=Nm(\mathfrak{p})\}$
@@
$\sum_{2\leq N(I)\leq X}Q^{2}(I)=\sum_{N(\mathfrak{p})\leq\sqrt{X}}|S_{2}(X,\mathfrak{p})|+O\left(\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}\right)$ (4.13)
Here, we observe that the sum on the right of (4.13) is exactly the same that we treated in the first half of the proof, without the conjugacy class condition on the prime ideal $\mathfrak{p}$. But the change in the consideration of the condition hardly changes anything in the estimation of the sum. We can start by using the estimate in (3.27) to see that for $Y\leq e^{(\log X)^{1-\delta}}$, for some $\delta>0$,
$\sum_{N(\mathfrak{p})\leq Y}|S_{2}(X,\mathfrak{p})|\ll\frac{X\log Y}{\log X}$ (4.14)
@@
$\sum_{Y<N(\mathfrak{p})\leq\sqrt{X}}|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\end{subarray}}\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)$ (4.15)
@@
$\hskip 56.9055pt+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\;\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{q})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)$ (4.16)
We observe that the only difference in between the equations (4.6) and (4.16) is the conjugacy class condition. Thus, we do a similar replacement of the inner sums with integrals, but this time, without the Chebotarev Density factor in front of the integrals. We make the following replacements:
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
@@
$E_{3}$ $:=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right)$
$E_{4}$ $:=\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right)$
@@
$|E_{3}|\ll\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\quad\text{and}\quad|E_{4}|\ll\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}$ (4.17)
Therefore, combining (4.14), (4.16) and (4.17), we have that
$\sum_{\begin{subarray}{c}N(\mathfrak{p})\leq\sqrt{X}\end{subarray}}|S_{2}(X,\mathfrak{p})|=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}$ $\int_{Y}^{N(\mathfrak{q})}\Psi^{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi^{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right)$ (4.18)

### v2

@@ 4 A New Formula for the Chebotarev Density in arbitrary finite Galois extensions
We will prove two theorems in this section, leading up to a new formula for the Chebotarev Density. We will be proving the results in a setting of a finite Galois extension $L/K$ such that $G=Gal(L/K)$ is the corresponding Galois group. We fix a conjugacy class $C\subset G$ which will be instrumental in our proofs. As mentioned above, the following results can be viewed as generalizations of the works of Alladi, Dawsey, Sweeting, Woo, Johnson, and the author himself. More precisely, the result that will be derived at the end of this section will prove to be a perfect generalization101010our result generalizes the choice of the Galois extension to equation (6.3) in [Se25].
Our first theorem in this section is a crucial one for our goal. It enables us to bring in the second order duality (Theorem 2.3) and the Chebotarev Density into the picture. This theorem can be treated as the backbone of this whole process to deduce the formula.
@@
$\sum_{2\leq N(I)\leq X}Q^{2}_{C}(I)=c_{K}\cdot\frac{|C|}{|G|}\cdot X+O\left(\frac{X(\log\log X)^{2}}{\log X}\right).$
@@
$Q_{C}^{2}(I)$ $:=\#\left\{\mathfrak{p}\subset\mathcal{O}_{K}:\;\mathfrak{p}\text{ is a prime; }I\subset\mathfrak{p};\;M_{2}(I)=N(\mathfrak{p})\;\text{and}\;\left[\frac{L/K}{\mathfrak{p}}\right]=C,\;\mathfrak{p}\;\text{is unramified}\right\}.$
By definition, it is clear that ideals in $\mathcal{O}_{K}$ contained in prime ideals only of a particular norm do not contribute to the sum in the LHS. As done in the proof of Lemma 3.4, in a similar fashion, we define the set $S_{2}(X,\mathfrak{p})$ as the set of ideals $I$ in $\mathcal{O}_{K}$ such that $N(I)\leq X$ with a unique and non-repeating prime ideal of the largest norm and $M_{2}(I)=N(\mathfrak{p})$. We note here that for any ideal $I\subset\mathfrak{p}$, for some prime $\mathfrak{p}$ such that $N(I)\leq X$ and $M_{2}(I)=N(\mathfrak{p})$, $N(\mathfrak{p})\leq\sqrt{X}$. Therefore, using Lemma 3.1, we have that
$\sum_{2\leq N(I)\leq X}Q_{C}^{2}(I)=\sum_{\begin{subarray}{c}N(\mathfrak{p})\leq\sqrt{X}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}|S_{2}(X,\mathfrak{p})|+O\left(\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}\right).$ (4.1)
@@
$|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}N(\mathfrak{M})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{M})\leq N(\mathfrak{p})\end{subarray}}\sum_{\begin{subarray}{c}N(\mathfrak{p})<N(\mathfrak{q})\\ N(\mathfrak{M.pq})\leq X\end{subarray}}1=\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}}\sum_{\begin{subarray}{c}N(\mathfrak{M})\leq\frac{X}{N(\mathfrak{pq})}\\ M_{1}(\mathfrak{M})\leq N(\mathfrak{p})\end{subarray}}1.$ (4.2)
@@
$|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{O}_{K}\\ \mathfrak{q}\text{ is a prime}\\ N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{pq})},N(\mathfrak{p})\right).$ (4.3)
@@
$\sum_{\begin{subarray}{c}N(\mathfrak{p})\leq Y\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}|S_{2}(X,\mathfrak{p})|\leq\sum_{N(\mathfrak{p})\leq Y}|S_{2}(X,\mathfrak{p})|\leq\sum_{N(\mathfrak{p})\leq Y}\Psi_{K,2}\left({X},\mathfrak{p}\right)\ll\frac{X\log Y}{\log X}.$ (4.4)
@@
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\;\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right).$ (4.5)
@@
$Y<N(\mathfrak{p}),\quad N(\mathfrak{p})<N(\mathfrak{q}),\quad N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{q})},\quad N(\mathfrak{p})\leq\sqrt{X}.$
@@
$\hskip 56.9055pt+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\;\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{q})}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right).$ (4.6)
To bring in the Chebotarev Density, we now interchange the inner sums of both the double sums in the RHS of (4.6) with suitable integrals. This is indeed one of the most crucial steps of the whole proof. We do the replacements as follows:
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\frac{|C|}{|G|}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t},$
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\frac{|C|}{|G|}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}.$
@@
$E_{1}$ $:=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\frac{|C|}{|G|}\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right),$
$E_{2}$ $:=\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\frac{|C|}{|G|}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right).$
@@
$=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\;\sum_{N(\mathfrak{M})\leq\frac{X}{YN(\mathfrak{q})}}\left|\sum_{\begin{subarray}{c}\max(Y,M_{1}(\mathfrak{M}))\leq N(\mathfrak{p})\leq\min\left(N(\mathfrak{q}),\frac{X}{N(\mathfrak{Mq})}\right)\\ \left[\frac{L/K}{\mathfrak{p}}\right]=C\end{subarray}}1-\frac{|C|}{|G|}\int_{\max(Y,M_{1}(\mathfrak{M}))}^{\min\left(N(\mathfrak{q}),\frac{X}{N(\mathfrak{Mq})}\right)}\frac{dt}{\log t}\right|.$ (4.8)
@@
$\ll\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}.$ (4.10)
Therefore, combining (4.7), (4.9), and (4.10), we have that
@@
$+$ $\frac{|C|}{|G|}\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right).$ (4.11)
@@
$+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{4}\sqrt{\log Y}}}\right).$ (4.12)
Now, we will evaluate the pseudo-integro-sums using further estimates. Surprisingly, the very same idea used above will work here too. Let us recall the definition of $Q^{2}(I)$:
$Q^{2}(I)$ $:=\#\{\mathfrak{p}\subset\mathcal{O}_{K}:\;I\subset\mathfrak{p};\;M_{2}(I)=Nm(\mathfrak{p})\}.$
@@
$\sum_{2\leq N(I)\leq X}Q^{2}(I)=\sum_{N(\mathfrak{p})\leq\sqrt{X}}|S_{2}(X,\mathfrak{p})|+O\left(\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}\right).$ (4.13)
Here, we observe that the sum on the right of (4.13) is exactly the same as that we treated in the first half of the proof, without the conjugacy class condition on the prime ideal $\mathfrak{p}$. But the change in the consideration of the condition hardly changes anything in the estimation of the sum. We can start by using the estimate in (3.27) to see that for $Y\leq e^{(\log X)^{1-\delta}}$, for some $\delta>0$,
$\sum_{N(\mathfrak{p})\leq Y}|S_{2}(X,\mathfrak{p})|\ll\frac{X\log Y}{\log X}.$ (4.14)
@@
$\sum_{Y<N(\mathfrak{p})\leq\sqrt{X}}|S_{2}(X,\mathfrak{p})|=\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\sqrt{X}\end{subarray}}\sum_{N(\mathfrak{p})<N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{p})}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right).$ (4.15)
@@
$\hskip 56.9055pt+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\;\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})\leq\frac{X}{N(\mathfrak{q})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right).$ (4.16)
We observe that the only difference between equations (4.6) and (4.16) is the conjugacy class condition. Thus, we do a similar replacement of the inner sums with integrals, but this time, without the Chebotarev Density factor in front of the integrals. We make the following replacements:
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t},$
$\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)\quad$ $\text{by}\quad\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}.$
@@
$E_{3}$ $:=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<N(\mathfrak{q})\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right),$
$E_{4}$ $:=\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\left(\sum_{\begin{subarray}{c}Y<N(\mathfrak{p})<\frac{X}{N(\mathfrak{q})}\end{subarray}}\Psi_{K}\left(\frac{X}{N(\mathfrak{p}\mathfrak{q})},N(\mathfrak{p})\right)-\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}\right).$
@@
$|E_{3}|\ll\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\quad\text{and}\quad|E_{4}|\ll\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}.$ (4.17)
Therefore, combining (4.14), (4.16), and (4.17), we have that
$\sum_{\begin{subarray}{c}N(\mathfrak{p})\leq\sqrt{X}\end{subarray}}|S_{2}(X,\mathfrak{p})|=\sum_{Y<N(\mathfrak{q})\leq\sqrt{X}}$ $\int_{Y}^{N(\mathfrak{q})}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}+\sum_{\sqrt{X}<N(\mathfrak{q})\leq\frac{X}{Y}}\int_{Y}^{\frac{X}{N(\mathfrak{q})}}\Psi_{K}\left(\frac{X}{tN(\mathfrak{q})},t\right)\frac{dt}{\log t}$
$+O\left(\frac{X\log Y}{\log X}\right)+O\left(\frac{X\log X\log\log X}{e^{c_{1}\sqrt{\log Y}}}\right).$ (4.18)